### v3

Abstract
Large language models(LLMs) have shown its outperforming ability on various tasks and question answering. However, LLMs require substantial memory storage on low-resource devices. More critically, the computational speed on these devices is also severely limited. In this paper, we propose SPA(Side Plugin Adaption), a lightweight architecture for fast on-devices inference on the constraints of strict on-devices computation and memory constraints. Compared with other on-devices seq2seq generation, SPA could make a fast and stable inference on low-resource constraints, allowing it to obtain cost effiency. Our method establish an interaction between a pretrained LLMs on-cloud and additive parameters on-devices, which could provide the knowledge on both pretrained LLMs and featured personal feature. Further more, SPA provides a framework to keep feature-base parameters on low computational devices while leave the parameters containing general information on the high computational devices.  [^1] Our code is public at https://github.com/OceannTwT/spa
Keywords:
Personalized LLM Cloud-device Collaboration Inference Acceleration.
1 Introduction
Large Language Models (LLMs)[4, 8, 33] have demonstrated incredibly powerful capabilities in supporting downstream tasks. These models perform exceptionally well across various downstream tasks. Additionally, different LLMs such as GPT-4[25], Llama[33], and ChatGLM[10, 38] exhibit similar abilities in handling tasks across multiple dimensions[2]. However, deploying and computing LLMs on resource-constrained devices introduces significant memory and computational pressure on these devices[1]. Therefore, researching reliable methods for deploying LLMs on-devices becomes critically urgent.
Figure 1: Three setting for model deployment.
To address the issue of deploying large language models (LLMs) on devices, various methods have been proposed to adapt them for edge devices. Subformer [29] reduces storage and computation requirements by sharing parameters. Similarly, EdgeFormer [12] proposes a balanced sharing scheme, exploring the impact of different architectures on device deployment. LLMCad [37] employs a collaborative model configuration to achieve lightweight deployment, where a small model generates simple words and a high-accuracy model verifies and corrects errors.
Additionally, research has proposed parameter-efficient fine-tuning methods for LLMs [14, 15, 18], utilizing adapter strategies to flexibly adapt to downstream tasks while preserving model capabilities. Adapters are stored separately from the base model and require only a few parameters to customize for downstream tasks [16]. For on-device deployment, adapters can be used for incremental deployment, integrating incremental knowledge into existing models to support downstream tasks and obtain user attributes [21]. However, this approach requires consideration of data transmission requirements between devices [31], and reducing transmission latency is a significant challenge in the design of separated models.
Our research proposes a method called Side Plugin Adaptation (SPA), which achieves the deployment of large models on-devices by separating adapters from pre-trained models. Specifically, our method introduces an side adapter classifier. This classifier allows us to choose between leveraging the inherent capabilities of the original model and integrating feature information generated by adapters. In this way, we aim to minimize the frequency of adapter usage while maximizing the utilization of common-sense knowledge inherent in the original large language model. This strategy offers a dual advantage: leveraging the unique feature attributes of adapters while maximizing the inherent capabilities of a widely based model. Compared to traditional on-devices model setups, our method significantly improves inference speed.
We comprehensively evaluate our approach on various datasets across different domains. These evaluations not only emphasize the advantages of our method compared to traditional approaches but also highlight its potential in improving overall model capability while achieving significant adapter effects. In addition to enhancing text generation and semantic understanding services, our approach demonstrates a stronger ability to understand textual features, better meeting user needs.
Our contributions can be summarized as follows:
  • $\bullet$
    We innovatively propose the SPA framework, deploying large models on-device via adapters. Our method significantly reduces the difficulty of fully deploying large models on the edge and presents a feasible and efficient solution.
  • $\bullet$
    We introduce a classifier to determine whether to use adapter content for inference. This innovative classifier design significantly improves the efficiency of cloud server to edge transmission while maintaining task performance.
  • $\bullet$
    SPA has been thoroughly tested across multiple datasets. Experiments show that our framework performs well on-devices. Additionally, SPA-LST can achieve results close to the original model with lower transmission latency.
2 Related Work
2.1 Parameter-Efficient Transfer Learning
In recent years, efficient parameter-efficient transfer learning has become a popular method for fine-tuning large language models. It achieves performance close to full fine-tuning by refining a subset of parameters. Among these methods, adapters, as introduced by [14], incorporate two additional feed-forward layers, enhancing their adaptability to downstream tasks. Similarly, LoRA [15] employs a rank reduction followed by rank promotion method on the original language model. In the context of language model fine-tuning, prefix tuning [19] involves introducing a trainable prefix vector before the keys and values in the multi-head attention mechanism, producing effects similar to the first two methods.
The MAM adapter [13] synthesizes these techniques and identifies superior results in parameter fine-tuning. Moreover, LST [32] provides a method for saving space during parameter fine-tuning. It achieves optimal parameter fine-tuning with the same memory consumption by leveraging adapters on the device. This approach significantly reduces memory usage when training large language models on edge devices. Meanwhile, the QLora[9] method combines quantization rules with LoRA’s rank promotion and demotion scheme, optimizing training resources and space savings.
Furthermore, the CuQA[18] method offers a classifier that selects between the original language model and edge models based on the relevance of training and testing data during training. Through this selection process, we observe that efficient parameter fine-tuning methods can effectively embed new knowledge, thereby expanding existing language models to incorporate new knowledge bases in support of specific tasks. Our method leverages the parameter-saving capabilities of efficient parameter-efficient transfer learning and the adaptive information contained in parameter fine-tuning inspired by CuQA. We provide robust support for various generative tasks, demonstrating enhanced capabilities.
2.2 On-device ML optimization
For on-device sequence-to-sequence models, various techniques have been adopted to reduce the parameters of pre-trained models, including methods such as quantization[20, 22] to reduce model size, knowledge distillation[27, 35] techniques that learn from large models and transfer knowledge to smaller models, and techniques such as compression[11, 26] and pruning[3, 6, 17] to reduce the number of model parameters. Some methods like Edgeformer [30] achieve model reduction by sharing model parameters. Low-rank decomposition is also used to reduce model size [5]. Methods like Delight [23] achieve model size reduction by compressing attention mechanisms. These techniques are adjusted and adapted based on the parameters of the Transformer [34] architecture, and they yield effective parameter reduction results for models with relatively fewer parameters.
Given that a considerable portion of these methods can only be applied to models with small model, they may not produce substantial optimizations in the context of large models. Computational and storage resources on devices may not be sufficient to accommodate LLMs. Therefore, we propose a framework for hybrid cloud server and on-device sequence-to-sequence model. The framework aims to meet the needs of certain downstream tasks and user characteristics by combining the capabilities of the cloud and edge.
3 Side Plugin Adaption
Given the computational and storage capabilities on the edge, our concerns can be broken down into the following points:
  • How much reduction in computation time and space cost can our model achieve?
  • Does its capability improve compared to the original model under these costs?
  • Whether our network latency matches expectations, and whether the overall latency significantly decreases compared to the original model’s cost.
Figure 2: Cloud-base and on-devices collaboration model: SPA.
These are some fundamental requirements and abilities for a small on-device model. Considering the need to learn user features and context-related information for text generation, we propose adding a classifier after the output of the transformer. We explore two settings for this classifier, comparing them to the original edge adapter’s results to assess downstream task support.
In the first setting, a binary choice approach is used, where the edge model is trained from fine-tuning and the output is generated either the origin LLM or side block results. The classifier selects the side whose output better aligns with the desired features, enabling the classifier to choose that path.
The second setting, called the plugin method, involves using the classifier to decide whether to combine the edge model with the base model stored in the cloud. This benefits by allowing the edge model to learn only the differences in features relevant to the current task. General information is then obtained by utilized the base model’s results, which reduced latency cost on edge devices resources.
Specifically, The function $f$ is the single layer on origin LLMs, and the function $g$ is the corresponding layer on side model. the methods above could conclude to these two functions:
$h=f(x)+\sigma g(x)$ (1)
$\sigma=\arg\max{\textbf{M}_{c}f(x)}$ (2)
We employed a linear layer of $H\times 2$ as our classifier. The advantage of this design lies in our ability to use the Transformer’s output as the hidden state, which is then softmax to a range between [00, $1$], thereby obtaining a specific distribution for classification choices $\sigma$. Through such a linear mapping, the model can learn whether predicting a certain token is better accomplished by selecting the existing model’s output (typically leveraging previously predicted information like "New + York"), or by employing a side-end incremental model (which relies more on user-specific feature information), thus determining the optimal choice. Considering $\sigma$ is learnable, the loss function in the whole model could be depicted as:
$\mathcal{L}_{LM}=-\sum_{i=1}^{T}\log_{(\Theta+\sigma\Delta\Theta)}{\textbf{P}(x_{i}|x_{<i})}$ (3)
Here we keep $f(x)$ frozen and train the side model $g(x)$ and classifier $\sigma$. Unlike a fixed classifier $\sigma^{\prime}$ for text similarity on the original model, our linear-layer-based classifier $\sigma$ can determine if a predicted word exists in the knowledge base of the base language model. This approach reduces inference times for highly repetitive word sequences, storing more general information in the cloud and generating user-specific features on the edge device. Compared to EdgeFormer, our model’s main bottleneck lies in the communication latency[31] between the edge and server. This can be managed by controlling communication frequency and times for an optimal value.
EdgeFormer[30] scales down the original model by factors like four or eight, sharing parameters to achieve a smaller model with similar capabilities and fewer FLOPs. However, this design demands greater computational and proficiency requirements for edge devices. While our methods set the edge model small, its prediction speed relative to total latency is quite small. In this case, the edge devices computational ability would not be a bottleneck for predicting the results on-devices.
For communication between cloud servers and edge models, we adopt the architecture used in LST. LST improves memory capacity from back propagate. For prediction, if our transformer has $L$ layers, Parallel adapter[13] or LoRA methods would cause a large cost on communication latency. Specifically, the latency including three parts. $T_{pretrained}$ indicates the original pretrained model computing latency. $T_{on-devices}$ indicates the edge devices model computing latency, it could be Adapter part, LoRA part or LST part for on-devices inference. What’s more, $T_{net}$ is the network latency when transmitting data between cloud and devices. Then, the total latency could be evaluated as:
$T_{total}=T_{on-devices}+T_{pretrained}+T_{net}$ (4)
While $T_{pretrained}$ keep the same for identical model. For model on edge devices, the inference latency is lay on the CPU or GPU computing capability, denoted as $f^{e}$. $F_{data}$ is the FLOPs required to compute on edges, $C_{devices}$ indicates the numbers of CPU or GPU are available to compute on this model. Then:
$T_{on-devices}=\frac{F_{data}\times{C_{devices}}}{f^{e}}$ (5)
For the network latency when transmitting data between cloud and devices. Assume the TCP/IP connection latency is $\tau$, The $T_{net}$ could be formulated as:
$T_{net}=\mathcal{M}\times(\tau+T_{data})$ (6)
Where $\mathcal{M}$ is the transmission times for one token prediction, it would be changed for different architecture, because we have to transfer back the hidden state from side model to base model. LST effectively merges the process in scenarios where edge-cloud connectivity is involved.
4 Experiments Setup
4.1 Datasets
We evaluated our approach on the following training dataset.
  1. 1.
    XSum[24]: There are 227, 000 summaries of online articles generated by the BBC available on XSum, and we assessed the relationships between the articles and their corresponding summaries.
  2. 2.
    CNN Daily Mail[7]: The CNN Daily Mail dataset consists of 312, 000 summaries of CNN Daily Mail articles and its origin articles.
  3. 3.
    CoQA[28]: CoQA datasets contains 127, 000 questions encompassing 8, 000 conversation-based question-answer pairs, constituting a large-scale QA system spanning various domains.
  4. 4.
    SciQ[36]: The SciQ dataset comprises 13, 679 publicly available scientific exam questions covering diverse disciplines such as biology, chemistry, physics, and others. Answers to specific questions are obtained through contextual information.
4.2 Implement Details
[…]
5 Experiments
The performance of SPA surpasses that of LST. Table 1 shows that, under the premise of lightweight parameters, the blocks trained by SPA can effectively guide the original model. Consequently, even with a smaller computational load at the edge, the model can exhibit strong inference capabilities and enhance performance for the given task.
Table 1: The Rouge-L scores on LST model, one-shot instruct and SPA model on various datasets.
Model XSum CNN-DM CoQA SciQ
LLaMA-7B + one-shot 17.32 22.36 15.32 17.21
LLaMA-7B + LST 28.18 32.15 31.24 23.24
LLaMA-7B + SPA 35.52 39.22 37.30 25.38
Table 2: Latency evaluated from Different efficient tuning methods. Measuring new parameters percentage rate, Inference latency, Network latency and Total latency on prediction for 50 new tokens. Ratios indicates the ratios of direct transmission times and pretrained model inferencing times. The Rouge-L scores on different PEFL methods and our model on various datasets.
Model % Param Infer. latency Net. latency Total latency Ratios XSum CoQA
LLaMA-7B + LoRA 0.19 3.26 6.37 9.63 32.0 38.24 40.32
LLaMA-7B + Adapter 0.38 3.32 12.56 15.88 64.0 38.18 40.51
LLaMA-7B + LST 0.2 3.29 0.31 3.60 1.0 28.78 31.24
LLaMA-7B + SPA 0.21 3.30 0.18 3.48 0.62 35.52 37.30
In the SPA architecture, network transmission is a crucial aspect, particularly in the context of SPA. Therefore, as the discussion in Section 3, it’s essential to compare the previous model with the SPA model and estimate the corresponding latency required for predicting the entire model. Given that our model is relatively large, we cannot deploy the entire model at the edge as in the case of Edgeformer. In the collaborative process between cloud-based and on-device components, transmitting critical information is necessary. Consequently, we will be conducting inference latency experiments to assess this aspect.
5.1 Latency on Prediction
To address concerns about network latency and overall inference delay, we conducted comparative experiments on data exchange for these models. Additionally, we compared the original Adapter, LoRA, and LST methods, evaluating them based on data exchange times and overall model latency. We performed these assessments using the same training configuration as mentioned above.
We also uniformly evaluated the time taken to generate 50 tokens. Additionally, our single-sided model and the baseline model were placed within a server cluster to simulate network transmission delays in an ideal network environment. This was done to assess the corresponding network transmission latency. The respective metric results obtained from the testing are recorded in Table 2. the difference between single and multiple communication latency is substantial, with inference costs relying mainly on cloud server computing power.
Table 3: The Rouge-L scores on different settings and SPA model on various datasets.
Model Ratios XSum CNN-DM CoQA SciQ
LLaMA-7B + S1 1.0 36.24 40.55 38.81 25.91
LLaMA-7B + S2 0.68 30.29 34.19 32.11 23.71
LLaMA-7B + S3 1.0 28.18 32.15 31.24 23.24
LLaMA-7B + SPA 0.62 35.52 39.22 37.30 25.38
Table 4: Different training data sizes on datasets.
Data sizes Xsum CNNDM CoQA SciQ
$\text{SPA}_{\text{small}}$ 10000 20000 4000 5000
$\text{SPA}_{\text{medium}}$ 100000 150000 40000 10000
$\text{SPA}_{\text{full}}$ 227000 312000 127000 13679
The network latency for the SPA model is superior compared with other methods. It’s noticeable that SPA exhibits favorable characteristics. It can employ the classifier to choose between a pre-trained model or a model with an integrated adapter for predictions. Due to the LST structure employed on-devices, it offers a more independent architecture compared to methods like LoRA and Adapters.
5.2 Impact of Side Model
We have assessed the impact of the edge model on the overall model and analyzed the word repetition rates between the edge model and the original model. Based on this analysis, we evaluate whether the results of our model meet our expected outcomes when selecting the edge model for inference. This metric can also be used to assess the extent of the difference between the new model and the original model and the proportionality with the edge model we have chosen. We show the evaluation of side model effect on Table 3.
We could find that while the origin cloud model and on-devices model generate different results, the classifier $\sigma$ would make a option that choose a more ideal information from the two alternations. When it comes to a same result, the classifier would choose to pick the cloud one for less network communication cost.
5.3 Analysis on Training Data Sizes
To assess the impact of training data sizes on model performance, we train the SPA model and utilize different sizes of training sets, as specified below:
We trained the models using the same network architecture and hyperparameters, with a consistent training epoch of 15 rounds. Text generation was performed using beam search.
Table 5: The Rouge-L scores on different training data sizes on various datasets.
Data sizes Xsum CNNDM CoQA SciQ
$\text{SPA}_{\text{small}}$ 31.5 36.2 34.7 22.1
$\text{SPA}_{\text{medium}}$ 33.8 38.4 36.2 23.8
$\text{SPA}_{\text{full}}$ 35.5 39.2 37.3 25.4
Table 6: Evaluations on side model ratios.
Data sizes $\text{SPA}_{\text{small}}$ $\text{SPA}_{\text{medium}}$ $\text{SPA}_{\text{full}}$
Usage Percentage 82.3 71.5 61.8
From the results, the model’s performance improves across multiple datasets with an increase in training data sizes, especially from small to medium data. This indicates that augmenting training data enhances the model’s generalization ability. However, as data sizes continues to increase to the full set, the performance improvement becomes less pronounced, indicating the presence of marginal effects. Considering the escalating computational resources, medium data size offers a favorable performance-to-cost ratio.
Additionally, we recorded the proportion of adapter usage by the model, as shown in Table 6. As data sizes increases, the percentage of adapter usage decreases, with a higher probability of selecting the original model’s output. This further validates that increasing data sizes enhances the model’s inherent language understanding and generation capabilities.
Increasing training data sizes significantly improves downstream task performance. However, there is a diminishing marginal effect with increased data sizes. Considering computational resources, a medium data size provides the best performance-to-cost ratio. More data reduces reliance on adapters, improving the model’s language understanding capabilities.
55551010101015151515202020202424242428282828323232323636363640404040Num of TrainningRouge-LXSumLSTParallel LST
55551010101015151515202020202525252529292929333333333737373741414141Num of TrainningRouge-LCNN-DMLSTParallel LST
Figure 3: Model Rouge-L scores comparison on parallel/sequential side model.
5.4 Different Setting on Side Plugin
SPA outperforms both LST and the pre-trained model, showcasing better expressive capabilities. In order to identify the source of this enhanced performance, we conducted ablation experiments.
We present three configurations in our study. The three configurations we marks from S1 to S3. In the first setup, referred to as "Full Integration, " the edge adapter directly interfaces with the original model. This arrangement allows the edge adapter to effectively learn and correct errors within the original model. In addition, the setup of S1 can be viewed as a parallelized LST setup, which has the advantage that he may perform better on generative tasks than the original serial option of LST.
Table 7: The Side Model Effect evaluation.
Model % Param M Diff rates Diff Pick rates
LLaMA-7B + SPA 0.21 0.62 78.2 56.8
In the second setup, a choice is made between the edge-adapted model and the original model. The classifier utilizes the original model’s output to make decisions between the results of the original model and the edge-adapted model. In this case, the edge on-devices model operates with a higher degree of independence from the original model.
The final setup involves the existing LST model, which can be viewed as relying entirely on the edge adapter for result generation. By comparing our model with these three setups, we aim to determine which configuration best aligns with the requirement for consistency between efficiency and capability that the model needs to achieve. The capabilities demonstrated by different experiments are summarized in Table 3.
Figure 4: The summary case study on classifier option, The pretrained model choices were marked as blue, while SPA model choices were marked as black.
The improvement of model performance in downstream tasks primarily relies on the adjustments made by the adapter to capture the incremental differences. As illustrate on Table 3, While the introduction of the classifier in the SPA model leads to a slight performance drop compared to the scenario where adapters are integrated in all cases, the performance reduction is relatively minimal. However, the improvement in inference speed is substantial. Considering our requirement for quick responses at the edge, such a classifier aligns better with our vision and needs. Simultaneously, for more fundamental information, predictions are made using the original model. The classifier significantly reduces the latency caused by continuous calls to the edge-side model during predictions, ensuring that the model learns both user and task features effectively.
We evaluate the performance of the sequence to sequence task between the original model and the model after parallel processing, and make corresponding comparisons under different datasets, and determine its performance under the corresponding task by the Rouge-L metrics.
In the parallelized design, the performance for our generative class of tasks is consistently higher than the original serialized design. As shown on Figure 3, such a parallelized design also highlights the plug-and-play nature of our model. After a very small number of rounds, the performance of parallelization is already more prominent than 20 rounds of generative training. Thus the use of a parallelized design in this paper is also in line with the corresponding sequence-to-sequence generation task.
5.5 Case Study
As shown on Figure 4. By detecting and processing the impact of classifiers on test cases, including the labels and branches used, we assess their influence and examine whether predictions are based on the original corpus or newly embedded knowledge. We analyze the top five results, observing key activation moments of the end-to-end model, as well as the predictions of the original pre-trained model at these moments.
The original model is primarily based on the usage of existing corpora and generated pronouns. These pronouns are generated within the original LLM, enabling it to learn the desired outcomes that users want to predict. For specific feature information, the output is primarily associated with embedded sentiment words and pronouns. These features are learned from the SPA model and are applicable to user profile features or downstream task requirements.
6 Conclusions
This paper introduces a model suitable for cloud-base and on-devices collaboration. This model leverages the capabilities of large models for sequence-to-sequence tasks and utilizes user-specific features from the edge model to guide the overall large model towards our desired objectives. Through a classifier, we can select the most suitable generated sentences. Our approach presents a new paradigm in addressing issues related to edge model deployment, providing valuable insights for future work. We believe that this approach is groundbreaking and can serve as a valuable reference for future research.